# Unpacking Positional Encoding in Transformers: A Spectral Analysis of Content-Position Coupling

### v1

@@ Abstract
Positional encoding (PE) is essential for enabling Transformers to model sequential structure. However, the mechanisms by which different PE schemes couple token content and positional information—and how these mechanisms influence model dynamics—remain theoretically underexplored. In this work, we present a unified framework that analyzes PE through the spectral properties of Toeplitz and related matrices derived from attention logits. We show that multiplicative content-position coupling—exemplified by Rotary Positional Encoding (RoPE) via a Hadamard product with a Toeplitz matrix—induces spectral contraction, which theoretically improves optimization stability and efficiency. Guided by this theory, we construct synthetic tasks that contrast content-position dependent and content-position independent settings, and evaluate a range of PE methods. Our experiments reveal strong alignment with theory: RoPE consistently outperforms other methods on position-sensitive tasks and induces “single-head deposit” patterns in early layers, indicating localized positional processing. Further analyses show that modifying the method and timing of PE coupling, such as MLA in Deepseek-V3, can effectively mitigate this concentration. These results establish explicit content-relative mixing with relative-position Toeplitz signals as a key principle for effective PE design and provide new insight into how positional structure is integrated in Transformer architectures.
@@ 1 Introduction
Transformer architectures have become the prevailing paradigm for sequence modeling across diverse domains owing to their powerful capacity to capture long-range dependencies. A cornerstone of their success is the attention mechanism’s ability to capture dependencies regardless of token distance. However, lacking inherent order sensitivity, Transformers critically rely on positional encoding (PE) to inject information about token positions and enable the modeling of sequential structure vaswani2017attention . The field has seen a proliferation of diverse PE variants, ranging from simple additive and learnable absolute embeddings to relative biases raffel2020exploring and sophisticated multiplicative schemes like Rotary Positional Encoding (RoPE) SU2024127063 . Yet, despite extensive empirical evaluation comparing their performance on downstream tasks, our fundamental understanding of the internal mechanisms by which these different PE variants specifically couple content and positional information, and how this coupling influences the model’s dynamics and internal learned structure, remains limited.
Existing evaluations of PE methods often rely on aggregate metrics such as perplexity HuaFoPE2025 or on length-extrapolation benchmarks kazemnejad2023impact . While providing valuable performance comparisons, these benchmarks shed little light on how positional signals are actually processed or on the mechanisms by which different encodings influence model behavior. Moreover, a focus on length-generalization often implies that, within a fixed sequence length, all PE schemes perform similarly. However, to our knowledge, no prior work has systematically confirmed or challenged this perception regarding within-length performance differences. This lack of understanding and empirical investigation highlights a critical gap in developing a unified theoretical framework that rigorously connects the specific design principles of PE (i.e., how they formulate positional coupling) to their impact on the attention dynamics and learned internal structure.
In this paper, we bridge this gap by introducing a spectral-theoretic framework complemented by targeted empirical studies. As illustrated in Figure 1, our method begins with the decomposition of each token embedding into content and position components ($c_{i}$ and $p_{i}$). We then assume that the position component gives rise to a Toeplitz-structured term in the attention logits, and we characterize how various PE schemes—whether additive biases or multiplicative rotations—map to distinct Toeplitz (or related) matrix operations within the attention computation.
Figure 1: Schematic Overview. (Left) Our framework analyzes token decomposition and how PE mechanisms (additive, e.g., adding Toeplitz $\mathbf{B}$; multiplicative, e.g., Hadamard with relative-position Toeplitz $G_{\mathbf{e}}$) structure attention logits. (Right) Multiplicative coupling theoretically induces spectral contraction (tighter eigenvalue bounds), predicting better learning/generalization and single-head deposit patterns. Additive mechanisms show smaller gains and more diffuse positional processing.
Our framework shows that content–position coupling can take two principal forms in the attention logits: (i) additive mechanisms, where position information (e.g., relative-position encodings) contributes by adding an explicit Toeplitz matrix $\mathbf{B}$ to the logits; and (ii) multiplicative schemes, most notably RoPE’s mechanism involving a Hadamard product with the relative-position Toeplitz matrix $G_{\mathbf{e}}$. Applying spectral analysis techniques, including Szegő’s theorem and Schur-product inequalities, we demonstrate theoretically that multiplicative coupling induces spectral contraction—yielding tighter eigenvalue bounds for the relevant matrix structures. This spectral contraction is known to improve optimization stability and efficiency. Consequently, schemes with this property are expected to learn the positional coupling function more rapidly and robustly in shallow layers, leading to its localization within a small set of attention heads. By contrast, additive mechanisms, lacking a direct content–relative position interaction in the logit, are predicted to underperform on tasks that critically depend on combining content with relative positional information.
To empirically validate our theoretical predictions, we design two synthetic tasks: Task 1, which requires content–relative positional reasoning, and Task 2, a position-agnostic control. We evaluate several PE methods (RoPE, Absolute, Relative (T5-style), ALiBi, NoPE) and conduct systematic layer- and head-wise ablations—zeroing out each head’s outputs—to probe the resulting internal structures. We also examine alternative coupling strategies, including MLA variants inspired by Deepseek-V3, to assess how different mechanisms and mixing timings shape positional processing.
Our results align closely with theory. On Task 1, requiring content-relative coupling, RoPE’s multiplicative mechanism consistently outperforms other methods, yielding superior generalization. Head-wise ablations confirm the predicted localization: RoPE produces an early-layer “single-head deposit,” wherein one shallow head carries the bulk of positional computations. By contrast, additive-bias schemes (Absolute, Relative, ALiBi) and NoPE exhibit diffuse or negligible head specialization in both tasks, and RoPE itself shows no deposit pattern on Task 2. Furthermore, MLA variants—specifically those employing a concatenation-based coupling mechanism (e.g., before the inner product)—produce a more distributed “average deposit” pattern. These findings provide strong empirical evidence that the form of content–position coupling fundamentally determines the model’s internal structure and learning dynamics.
  • We introduce a unified spectral-theoretic framework analyzing positional *interaction* in attention logits as structured (Toeplitz) matrices, distinguishing fundamental additive (bias/summation) and multiplicative (Hadamard) coupling mechanisms.
  • We show theoretically that multiplicative coupling (e.g., RoPE) induces spectral contraction—tighter eigenvalue bounds—which theory links to improved optimization stability. We empirically validate that PE methods employing these fundamental mechanisms exhibit predicted performance differences and deposit patterns (e.g., RoPE’s shallow single-head localization) on tasks requiring content-position coupling.
  • Via systematic layer- and head-wise ablations, we further investigate alternative strategies beyond fundamental mechanisms, such as MLA’s concatenation-based design. We demonstrate that these approaches yield contrasting, more distributed deposit patterns (e.g., average deposit vs. single-head localization), providing insights into how processing and combination choices influence learned internal structures.
The remainder of this paper is organized as follows. Section 3 details our spectral-theoretic framework. Section 4 describes the experimental setup and main results. Section 5 presents the ablation studies and mechanistic analysis.
@@ 2 Related Work
Positional Encoding in Transformers. Positional encoding was introduced alongside the Transformer architecture to endow self-attention with sequence order information vaswani2017attention . Early approaches employed learnable positional embeddings shaw2018self , while subsequent work integrated position biases directly into the attention logits, giving rise to relative position encoding and its extensions press2022alibi ; li2024fire ; chi-etal-2023-dissecting ; chi2022kerple ; raffel2020exploring ; SU2024127063 ; zhao2025drope . Among these, Rotary Positional Encoding (RoPE) SU2024127063 stands out by applying a multiplicative bias to the $QK^{\top}$ matrix, and has been widely adopted in modern open-source models. Related efforts on sliding-window and length-generalization designs are also often framed as positional encoding variants kiyono-etal-2021-shape ; chen2024clex ; chen2024longlora ; ding2024longrope ; peng2024yarn . Recent RoPE-based improvements include Frequency-Partitioned Encoding (FoPE) HuaFoPE2025 and the Multi-head Latent Attention (MLA) mechanism in Deepseek-V3 liu2024deepseek .
@@
Mechanistic Analyses of Positional Encoding. Empirical studies have probed how Transformers internalize positional signals. Some works demonstrate that even without explicit positional embeddings (NoPE), models can learn positional distinctions wang-etal-2024-length ; zuo-etal-2025-position ; haviv-etal-2022-transformer ; barbero2024round , occasionally outperforming PE-equipped counterparts on select tasks kazemnejad2023impact . Explanations range from implicit causal masking zuo-etal-2025-position to dataset-dependent effects barbero2024round . Other analyses document characteristic behaviors of RoPE, such as “massive value” phenomena and rotation artifacts jinMassiveValuesSelfAttention2025 ; jonasson2025rotary .
@@
Toeplitz Structure and Spectral Theory. The action of positional encoding in attention can be unified under Toeplitz-matrix representations bottcher2000analysis ; grenander2002toeplitz ; gray2006toeplitz ; oppenheim1999discrete . Classical results on Toeplitz spectral distributions, including Szegő’s theorem, inform our theoretical framework. Moreover, the impact of matrix spectra on optimization convergence is well studied in numerical linear algebra and convex optimization contexts horn1985matrix ; horn1991topics ; boyd2004convex ; nesterov2004introductory .
3 Theoretical Structure
3.1 Preliminaries and Assumptions
In Transformer architectures, the attention mechanism relies on the inner product between the query vector $q_{i}$ and the key vector $k_{j}$ at positions $i$ and $j$, respectively. This inner product, $q_{i}^{\top}k_{j}$, serves as the foundation for computing attention scores. Different positional encoding (PE) strategies influence this computation by injecting positional information in distinct ways.
Absolute Positional Encoding vaswani2017attention introduces position information at the input layer by directly adding a position vector $p_{i}$ to each token embedding $e_{i}$:
$h_{i}=e_{i}+p_{i}.$ (1)
Consequently, the attention scores are computed based on position-infused vectors:
$\alpha_{i,j}\propto\exp\bigl{(}(e_{i}+p_{i})^{\top}(e_{j}+p_{j})\bigr{)}.$ (2)
The residual connections throughout the Transformer layers ensure that this form of position-content coupling persists at deeper levels.
Relative Positional Encoding raffel2020exploring modifies the attention logits post inner-product, by adding a relative position-dependent bias term $b_{j-i}$:
$\alpha_{i,j}\propto\exp\bigl{(}q_{i}^{\top}k_{j}+b_{j-i}\bigr{)}.$ (3)
Here, $b_{j-i}$ is often a learnable parameter, either directly or as a function of relative distance $j-i$ parameterized through shared weights. This method allows position-dependent modifications without altering the token representations themselves.
Rotary Positional Encoding (RoPE) SU2024127063 instead applies a position-specific rotation matrix $R_{i}$ (or $R_{j}$) to each query and key vector before the inner product is computed. Due to the properties of these rotational transformations, the resulting dot product admits a key equivalence:
$q_{i}\leftarrow R_{i}q_{i},\quad k_{j}\leftarrow R_{j}k_{j}\quad\Longrightarrow\quad(R_{i}q_{i})^{\top}(R_{j}k_{j})=q_{i}^{\top}R_{j-i}k_{j}.$ (4)
Here, $R_{j-i}$ is a rotation matrix determined solely by the relative position difference. RoPE thus achieves a tightly coupled interaction between content and relative position within the attention score computation.
Notably, even in the absence of explicit positional encoding—commonly referred to as NoPE haviv-etal-2022-transformer —large Transformer models have demonstrated an ability to implicitly encode position through learned patterns in attention heads or depth-wise representation dynamics. This motivates our first key assumption:
@@ 3.1 Preliminaries and Assumptions
Each token representation $x_{i}$ at position $i$ can be decomposed into a position-independent content component $c_{i}$ and a position-dependent component $p_{i}$, such that
$x_{i}=c_{i}+p_{i}.$ (5)
This decomposition provides a conceptual foundation to analyze how different PE mechanisms construct or interact with the positional component $p_{i}$, and how this interaction shapes attention.
To formalize the structural implications of positional interaction in attention, we invoke a classic object in matrix analysis:
@@
A matrix $A$ is a Toeplitz matrix if its entries are constant along each diagonal for some sequence ${a_{k}}$. An $N\times N$ Toeplitz matrix $T$ can be expressed as:
$T=\begin{bmatrix}a_{0}&a_{-1}&a_{-2}&\cdots&a_{-(N-1)}\\ a_{1}&a_{0}&a_{-1}&\cdots&a_{-(N-2)}\\ a_{2}&a_{1}&a_{0}&\cdots&a_{-(N-3)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{N-1}&a_{N-2}&a_{N-3}&\cdots&a_{0}\end{bmatrix},$ (6)
We denote this compactly as $T=[a_{i-j}]$.
Toeplitz matrices naturally emerge in systems with translation-invariant or shift-dependent structure, making them particularly suited to model attention patterns influenced by relative position.
Our second foundational assumption connects this matrix structure to positional information within the attention mechanism:
@@
The attention score matrix induced solely by the position-dependent components $p_{i}$ exhibits a Toeplitz structure:
$\mathbf{Attention}(p_{i},p_{j})=[a_{i-j}].$ (7)
@@
The equal sign is used here because the normalization operation does not change the Toeplitz property of the matrix.

### v2

# Deconstructing Positional Information: From Attention Logits to Training Biases

@@ Abstract
Positional encodings enable Transformers to incorporate sequential information, yet their theoretical understanding remains limited to two properties: distance attenuation and translation invariance. Because natural language lacks purely positional data, the interplay between positional and semantic information is still underexplored. We address this gap by deconstructing the attention-logit computation and providing a structured analysis of positional encodings, categorizing them into additive and multiplicative forms. The differing properties of these forms lead to distinct mechanisms for capturing positional information. To probe this difference, we design a synthetic task that explicitly requires strong integration of positional and semantic cues. As predicted, multiplicative encodings achieve a clear performance advantage on this task. Moreover, our evaluation reveals a hidden training bias: an information aggregation effect in shallow layers that we term the single-head deposit pattern. Through ablation studies and theoretical analysis, we proved that this phenomenon is inherent in multiplicative encodings. These findings deepen the understanding of positional encodings and call for further study of their training dynamics. Our code is available at https://github.com/Qihuai27/Deposit-Pattern-Research.
@@ 1 Introduction
Positional Encoding (PE) (Vaswani et al., 2017; Su et al., 2024) endows Transformer architectures with an understanding of sequence order, compensating for the permutation-invariant nature of self-attention (Vaswani et al., 2017). This component is critical to their success in both language (Vaswani et al., 2017) and vision (Dosovitskiy et al., 2021) tasks. The design of PE has evolved substantially, from the additive sinusoidal embeddings (Vaswani et al., 2017) of the original Transformer to trainable absolute embeddings (Shaw et al., 2018) and more sophisticated relative schemes like T5 biases (Raffel et al., 2020) and Rotary Positional Encoding (RoPE) (Su et al., 2024). This progression reflects a sustained effort to enhance generalization, extend context length, and improve model efficiency.
Different positional encodings model distance decay and translation invariance in different ways, but these differences are difficult to observe in the face of complex data with human grammatical rules. Current evaluation paradigms, which rely on aggregate metrics such as perplexity (Ermo,Hua et al., 2025) or length-extrapolation benchmarks (Kazemnejad et al., 2023), document these performance differences but fail to illuminate the underlying mechanisms of how positional signals are processed (Peng et al., 2024; Ermo,Hua et al., 2025). Therefore, theoretical research on positional encoding relies on synthesis tasks (Zuo et al., 2025; Kazemnejad et al., 2023), but some unexplained phenomena have emerged. For example, despite its strong theoretical properties, such as the attenuation characteristic that supports length generalization, RoPE can underperform simpler relative PEs or even models with no PE on certain tasks (Kazemnejad et al., 2023). This discrepancy highlights a critical gap in our understanding: How, precisely, do different PE schemes mediate the interaction between token content and position? This lack of mechanistic clarity hinders both the explanation of existing performance trade-offs and the principled design of future PE schemes.
To bridge this gap, our approach is grounded in a core principle of positional encoding: the relationship between any two tokens should depend on their relative distance, not their absolute positions. This property of translation invariance is formally captured by the structure of a Toeplitz matrix, in which all elements along any given diagonal are identical. This insight allows us to deconstruct the attention logit calculation and propose a unified, Toeplitz-based framework for analyzing PE mechanisms (Figure 1). In our framework, each token’s embedding is composed of a content component (what it is) and a position component (where it is). The central idea is that these position components collectively induce a Toeplitz structure in the final attention scores.
Figure 1: Schematic Overview. (Left) Each token is viewed as $x_{i}=c_{i}+p_{i}$. After the $W_{Q},W_{K}$ projections, their interactions form the four inner-product terms of the logit matrix. The framework unifies PE methods by how they introduce Toeplitz (translation-invariant) structure: additive PEs (Absolute, T5, ALiBi) contribute a Toeplitz bias $B$ or a Toeplitz $G_{q^{p},k^{p}}$; RoPE applies a relative-position Toeplitz kernel $G_{e}$ multiplicatively through a Hadamard product. (Right) These two modes produce distinct behaviors: RoPE’s multiplicative coupling gives faster learning and smaller generalization gaps, yet concentrates shallow-layer positional information into a “deposit head”.
This framework clearly distinguishes between two primary classes of PE. Additive methods, such as relative position biases, operate by adding a static Toeplitz matrix of position scores directly to the attention logits. In contrast, multiplicative methods like RoPE dynamically integrate positional information into the query and key representations. This creates a powerful content–position interaction, where the positional influence on attention scores is conditioned on the token’s content. While this coupling provides a significant advantage on tasks where meaning is tightly bound to relative location, we argue that it also introduces a strong inductive bias. This bias can manifest as a drawback, concentrating the learning of positional logic into a small, specialized subset of attention heads.
Observations based on the theoretical framework directly triggered our experimental design: we conduct experiments on a series of carefully designed synthetic tasks. We generate random sequences from a small vocabulary and use labeled “anchor” tokens to define two contrasting objectives: a position-sensitive task requiring reasoning about the relative positions of anchors, and a position-agnostic task dependent only on token counts. This controlled setup enables us to precisely isolate and evaluate the model’s positional reasoning capabilities. Our experiments confirm that RoPE decisively outperforms other methods on the position-sensitive task while underperforming on the position-agnostic one. More importantly, this setting reveals a striking artifact unique to RoPE: the “single-head deposit pattern”, where nearly all positional processing in the shallow layers becomes concentrated into a single attention head.
Further investigation confirms this phenomenon is unique to RoPE on tasks with strong content-position correlation. Through rigorous ablation studies and theoretical derivation—such as injecting absolute positional encodings or selectively applying RoPE to a subset of heads—we demonstrate that the deposit pattern is an intrinsic property of RoPE’s multiplicative structure, not an incidental training artifact. This finding also explains why Transformers using RoPE tend not to form the latent positional representations observed in models with additive or no PE (Zuo et al., 2025; Kazemnejad et al., 2023). We posit that this intense, structurally-induced specialization is a primary cause of the gap between RoPE’s theoretical promise and its practical performance.
Our contributions can be summarized as follows:
  • We propose a unified analytical framework using Toeplitz matrices that categorizes positional encodings into additive and multiplicative mechanisms, clarifying their distinct effects on attention logits.
  • We empirically identify and analyze RoPE’s performance paradox through targeted synthetic tasks, revealing a unique ‘single-head deposit pattern’ where positional logic becomes highly concentrated in shallow layers.
  • We conduct a causal analysis to demonstrate that the deposit pattern is an intrinsic property of RoPE’s multiplicative architecture, offering a mechanistic explanation for its observed performance paradox.
@@ 2 Related Work
Positional Encoding in Transformers. Positional encoding was introduced alongside the Transformer architecture to endow self-attention with sequence order information (Vaswani et al., 2017). Early approaches employed learnable positional embeddings (Shaw et al., 2018), while subsequent work integrated position biases directly into the attention logits, giving rise to relative position encoding and its extensions (Press et al., 2022; Li et al., 2024; Chi et al., 2023; 2022; Raffel et al., 2020; Su et al., 2024; Zhao et al., 2025). Among these, Rotary Positional Encoding (RoPE) (Su et al., 2024) stands out by applying a multiplicative bias to the $QK^{\top}$ matrix, and has been widely adopted in modern open‐source models. Related efforts on sliding‐window and length‐generalization designs are also often framed as positional encoding variants (Kiyono et al., 2021; Chen et al., 2024a; b; Ding et al., 2024; Peng et al., 2024). Recent RoPE‐based improvements include Frequency‐Partitioned Encoding (FoPE) (Ermo,Hua et al., 2025) and the Multi‐head Latent Attention (MLA) mechanism in Deepseek‐V3 (Liu et al., 2024).
@@
Mechanistic Analyses of Positional Encoding. Empirical studies have probed how Transformers internalize positional signals. Some works demonstrate that even without explicit positional embeddings (NoPE), models can learn positional distinctions (Wang et al., 2024; Zuo et al., 2025; Haviv et al., 2022; Barbero et al., 2024), occasionally outperforming PE‐equipped counterparts on select tasks (Kazemnejad et al., 2023). Explanations range from implicit causal masking (Zuo et al., 2025) to dataset‐dependent effects (Barbero et al., 2024). Other analyses document characteristic behaviors of RoPE, such as “massive value” phenomena and rotation artifacts (Jin et al., 2025; Jonasson, 2025).
@@
Toeplitz Structure and Spectral Theory. The action of positional encoding in attention can be unified under Toeplitz‐matrix representations (Böttcher and Silbermann, 2006; Grenander, 1958; Gray, 2006; Oppenheim and Schafer, 1999). Classical results on Toeplitz spectral distributions, including Szegő’s theorem, inform our theoretical framework. Moreover, the impact of matrix spectra on optimization convergence is well studied in numerical linear algebra and convex optimization contexts (Horn and Johnson, 2012; 1994; Boyd and Vandenberghe, 2004; Nesterov, 2013).
3 Methodology
3.1 Preliminaries: Positional Encoding in Attention
The Transformer attention score between a query $q_{i}$ at position $i$ and a key $k_{j}$ at position $j$ is based on their inner product, $q_{i}^{\top}k_{j}$. Since this computation does not depend on sequence order, Positional Encoding (PE) schemes introduce positional information by modifying either the token representations or the attention logits.
Absolute PE (Vaswani et al., 2017) adds a position vector $p_{i}$ to the initial embedding $e_{i}$:
$h_{i}=e_{i}+p_{i}.$
The query and key vectors derived from $h_{i}$ therefore mix content and position inside the representation, making the attention score depend on the absolute token positions.
Relative PE (Raffel et al., 2020) incorporates positional information directly into the attention logits through a learnable distance-dependent term:
$\alpha_{i,j}\propto\exp\bigl(q_{i}^{\top}k_{j}+b_{j-i}\bigr).$
A defining feature of relative PE is translation invariance: the positional contribution depends only on the displacement $j-i$, not on the absolute locations of $i$ and $j$. Within this family, T5 (Raffel et al., 2020) uses a learned table indexed by $(j-i)$, while ALiBi (Press et al., 2022) applies a fixed slope that increases linearly with distance. Thus both are translation-invariant, but differ in whether the distance mapping is learned (T5) or predetermined (ALiBi).
Rotary PE (RoPE) (Su et al., 2024) introduces position by rotating the query and key vectors with a position-dependent matrix $R_{i}$. A key identity shows that the resulting inner product depends only on the relative displacement:
$(R_{i}q_{i})^{\top}(R_{j}k_{j})=q_{i}^{\top}R_{j-i}^{\top}k_{j}.$
Thus RoPE implements positional encoding through a structured rotation that couples content and position via relative position, rather than through additive shifts or explicit bias terms.
3.2 A Unified Framework via Toeplitz Matrices
To analyze these mechanisms within a common framework, we introduce two foundational assumptions. The first is a conceptual decomposition of token representations.
@@ 3.2 A Unified Framework via Toeplitz Matrices
Any token representation $x_{i}$ can be conceptually decomposed into a position-independent content component $c_{i}$ and a position-dependent position component $p_{i}$, such that $x_{i}=c_{i}+p_{i}$.
Remark 3.1.
This decomposition is not a modeling constraint but a structural lens. It is supported in four complementary ways: (i) Design: Absolute PE explicitly defines position-dependent vectors $p_{i}$ (Vaswani et al., 2017); (ii) Emergence: Architectures lacking PE still develop internal positional directions (Zuo et al., 2025); (iii) Functionality: Ablating $p_{i}$ hurts performance more than ablating random vectors of equal norm (Section 5.4); (iv) Generality: We treat the removal of $p_{i}$ as a manifold-based coordinate subtraction rather than a rigid linear projection, ensuring compatibility with diverse PE schemes.
This allows us to analyze how PE schemes construct or interact with the positional component $p_{i}$. The core principle uniting relative PE schemes is translation invariance: the positional component of the attention score depends only on the relative distance $j-i$. This principle is perfectly embodied by a specific matrix structure.
@@
A matrix $\mathbf{T}$ is a Toeplitz matrix if its entries are constant along each diagonal, i.e., $T_{i,j}=a_{i-j}$ for some sequence $\{a_{k}\}$. It is defined by $2N-1$ unique values for an $N\times N$ matrix.
Such matrices naturally arise whenever a quantity depends only on the relative displacement $j-i$, making them the canonical representation of translation-invariant positional interactions. This leads to our second assumption, which connects the positional components of tokens to the structure of the attention matrix.
@@
Since the positional contribution to attention depends only on the relative displacement $j-i$, the Gram matrix formed by the position-dependent components naturally takes a Toeplitz form.
@@
Assumption 3.2 can be viewed less as a modeling assumption and more as a working hypothesis that captures the intended translation-invariant decay, and that the $p_{i}$ decomposed from Assumption 3.1 should have such a property. This assumption could be checked by sinusoidal absolute PEs and is explicitly enforced by construction in relative PE biases. It provides a powerful analytical tool for understanding the structure of positional information in the attention matrix.
Under Assumptions 3.13.2, the attention logits can be decomposed into interpretable content and position interactions. Let $q_{i}^{c},k_{j}^{c}$ be query/key vectors derived from content components and $q_{i}^{p},k_{j}^{p}$ be those from position components. Let $G_{v,w}$ denote the Gram matrix where $(G_{v,w})_{i,j}=v_{i}^{\top}w_{j}$.
For additive mechanisms (e.g., Absolute or Relative PE), the logit matrix is a sum of components:
$\displaystyle\mathbf{L}_{\text{Additive}}=G_{q^{c},k^{c}}+G_{q^{c},k^{p}}+G_{q^{p},k^{c}}+G_{q^{p},k^{p}}+\mathbf{B}$ (1)
Here, $\mathbf{B}$ denotes the explicit relative-bias matrix used only by relative encodings such as T5 and ALiBi. For absolute or learned absolute PEs, we instead have $\mathbf{B}=0$, and the Toeplitz positional structure is provided solely by the term $G_{q^{p},k^{p}}$. Thus, in all additive mechanisms the Toeplitz component arises either from $G_{q^{p},k^{p}}$ (absolute PE) or from $\mathbf{B}$ (relative PE), while the cross-terms $G_{q^{c},k^{p}}+G_{q^{p},k^{c}}$ remain the only channel through which content can interact with relative position.
In contrast, multiplicative mechanisms like RoPE induce a different structure. Following its formulation using complex numbers in  Su et al. (2024), the logit matrix becomes:
$\displaystyle\mathbf{L}_{\text{RoPE}}=\operatorname{Re}\left\{\left(G_{q^{c},k^{c}}+G_{q^{c},k^{p}}+G_{q^{p},k^{c}}+G_{q^{p},k^{p}}\right)\circ G_{\mathbf{e}}\right\}$ (2)
where $\circ$ is the Hadamard product and $G_{\mathbf{e}}$ is a Toeplitz matrix.